OpenStudy (anonymous):

How do you simplify 3x+13x-33/x^2-11x

OpenStudy (australopithecus):

well I would first split the fraction up using the rule (x+y)/dc = x/dc + y/dc

OpenStudy (australopithecus):

then factor

OpenStudy (australopithecus):

you follow?

OpenStudy (anonymous):

I understand what your saying, but how do I split it?

OpenStudy (australopithecus):

You can't split denominators though remember. A common mistake is xy/(d+c) =/= xy/d + xy/c (You cannot do this!!!!)

OpenStudy (australopithecus):

(3x+13x-33)/x^2-11x = 3x/(x^(2)-11x) + 13x/(x^(2)-11x) - 33/(x^(2)-11x)

OpenStudy (australopithecus):

its just like adding or subtracting fractions with the same denominator but in reverse :)

OpenStudy (australopithecus):

now you can simply factor the denominator and you will be able to simplify the fractions further

OpenStudy (australopithecus):

do you follow?

OpenStudy (anonymous):

not really :/

OpenStudy (australopithecus):

Do you see why 3x/(x^(2)-11x) + 13x/(x^(2)-11x) - 33/(x^(2)-11x) = 3x+13x-33/x^2-11x

OpenStudy (australopithecus):

I'm assuming your problem is